OpenStudy (babynini):

Convert polar eq. to rectangular cos2(Theta)=-1

OpenStudy (babynini):

@zepdrix

OpenStudy (babynini):

@sdfgsdfgs

OpenStudy (babynini):

@rvc

OpenStudy (babynini):

@perl

zepdrix (zepdrix):

Is that a square on the cosine? :)

OpenStudy (babynini):

Nope, it's\[\cos2(\theta)=-1\]

zepdrix (zepdrix):

Recall your Double Angle Formula for Cosine? :)

OpenStudy (babynini):

\[\cos^2(\theta)-\sin^2(\theta)=-1\]

zepdrix (zepdrix):

Wait, we're going to rectangular? 0_o but no r? Hmm interesting lol

OpenStudy (babynini):

yeah that's what's weird! xD I've tried multiplying r in but it gets pretty sketchy.

zepdrix (zepdrix):

How bout we... multiply both by r^2

zepdrix (zepdrix):

\[\Large\rm (r \sin \theta)^2-(r \cos \theta)^2=1\]I combined a few steps into one here, ok with this? :o

zepdrix (zepdrix):

Crappp my game is starting >.<

OpenStudy (babynini):

ai, do you need to go?

zepdrix (zepdrix):

Woops typo,\[\Large\rm (r \sin \theta)^2-(r \cos \theta)^2=r^2\]

zepdrix (zepdrix):

Yah :c Do you see how I got here? Can you continue from there maybe? :3

OpenStudy (babynini):

Hm, I'm not sure. But i'll work it out. You can go! thanks :)

zepdrix (zepdrix):

ok ill come back later if needed c:

OpenStudy (babynini):

okies. but the r on the right side would be negative, yeah?

OpenStudy (babynini):

@alekos

OpenStudy (perl):

This is the question $$\Large \cos(2\theta) = -1$$

OpenStudy (babynini):

yep

OpenStudy (perl):

We can use trig substitution $$ \Large{ \cos(2\theta) = -1 \\ 2\cos^2\theta -1 = -1 } $$

OpenStudy (babynini):

Okay, sounds good..

OpenStudy (babynini):

\[2\cos^2(\theta)=0\] simplify?

OpenStudy (perl):

correct

OpenStudy (perl):

We can use trig substitution $$ \Large{ \cos(2\theta) = -1 \\ 2\cos^2\theta -1 = -1 \\ 2\cos^2\theta = 0 \\ \cos^2\theta = 0 \\ \cos\theta = 0 \\ \theta = \arccos(0) \\ \theta = \frac{\pi }{2} } $$ now it is true that in polar to rectangular theta = arctan(y/x)

OpenStudy (perl):

actually that is the y axis, if you graph it

OpenStudy (perl):

so the rectangular equation is x = 0

OpenStudy (babynini):

oo yes this makes sense.

OpenStudy (perl):